4. A gold miner has a claim in the Yukon. Every day he goes to his claim and mines for gold. Assume that on any given day the probability that he will find gold is 0.005. Assume that his chance of finding gold on a particular day is independent of his finding gold on any other day.

a. what is the expected number of days that it will take him to find gold for the first time?

b. if he gives this experiment 2 years, what is the probability that he never finds gold?
Answer #1

this is geometric distribution with parameter p=0.005

a)

expected number of days that it will take him to find gold for the first time =1/p=1/0.005=200

b)

P( never finds gold in 2years (365*2=730 days)) =(1-0.005)730 =0.025754